# No-exclaves percolation on random networks

### v1

@@ Abstract
No-exclaves percolation (NExP) is a nonlocal percolation process in which the components are formed not only by the connected occupied nodes but also by the agglomeration of empty nodes completely surrounded by the occupied nodes. It has been studied in low dimensions, displaying such novel phenomena as the discontinuous transition to complete percolation. However, its characteristics in complex networks are still unexplored. In this paper, we study the NExP on random networks by developing mean-field solutions using the generating function formalism. Our theory allows us to determine the size of the giant no-exclaves component as well as the percolation threshold, which are in excellent agreement with Monte Carlo simulations on random networks and some real-world networks. We show that on random networks NExP exhibits three phases and two transitions between them: the phases are characterized by the presence or absence of not only the giant no-exclaves component but also the giant unoccupied component, which is the giant connected component composed solely of unoccupied nodes. This work offers theoretical understanding on the anatomy of phase transitions in the NExP process.
@@ 1 Introduction
Understanding percolation of complex networks has an impact on our ability to predict and control the behaviors of complex systems as it can provide essential information regarding the connectivity and robustness of complex systems stauffer ; error_attack ; newman_book ; dlee . Classical percolation models are based on the notion of connectivity in which a pair of nodes are considered a part of the same connected component if there exists at least one connected path between them stauffer . This model has provided a simple yet powerful theoretical framework across various phenomena, including epidemic spreading newman02 ; pastor02 , traffic flows transport , the stability of power grid power ; yang ; chilean , and the resilience of networks error_attack ; callaway ; cohen2000 . There have also been numerous studies on variants of the ordinary percolation models on networks that have modified the concept of connectivity such as $k$-core percolation kcore , bootstrap percolation bootstrap , $k$-selective percolation kim_goh , and mutual percolation buldyrev ; son ; baxter12 ; min14 .
@@
One of the most important applications of percolation theory in complex networks is the robustness of networks against random failures or attacks error_attack ; callaway ; cohen2000 ; cohen2001 ; holme . The robustness of networked systems is often measured by how the size of the largest connected component responds when some nodes in networks are failed error_attack ; holme . The premise underlying the measurement is that it is essential for nodes to maintain their connectivity in order to function properly error_attack ; buldyrev . Additionally, when nodes become non-functioning due to failure or detachment from the largest component these nodes are assumed to be unable to recover permanently in most studies on the robustness of networks error_attack ; newman_book . In reality, however, damaged nodes can often recover with communications to properly functioning nodes that surround the damaged nodes min14 ; recovery ; shekhman16 ; shang ; wu2022 . For instance, temporary breakdown in real-world networks such the brain and financial systems can recover as long as the damage is localized.
@@
The no-exclaves percolation (NExP) is a model of incorporating such a recovery rule that components of failed nodes that are completely surrounded by active (unfailed) nodes recover to be unfailed and merge into the surrounding unfailed component nexp . The failed components that are surrounded by unfailed nodes are referred to as “exclaves” which is the terminology in political geography and means a part of a district geographically isolated from the mainland by surrounding alien territories. Our model does not allow for exclaves, hence the name “no-exclaves percolation,” or NExP for short. After applying the no-exclaves rule for the recovery, we identify connected components and assess the robustness of networks using the concept of the giant NExP component. The no-exclaves rule in our model is closely related to the no-enclaves percolation on two dimensions offering a theoretical explanation for a motor-driven collapse of cytoskeletal systems sheinman ; gunnar ; reply ; hu2016 . Note that the no-enclaves rule does not readily apply to complex networks; the no-exclaves rule, however, is applicable to generic networks, while keeping the ingredient of nonlocality intact. Furthermore, the no-exclaves rule implies long distance communications between active nodes in some sense castellano2020 . From this perspective, it might have potential applicability to quantum communication and quantum network issues as well meng ; extended .
@@
It is known from numerical simulations that in low-dimensional systems the no-exclaves rule significantly affects the behavior of percolation transitions in spatial networks and Euclidean lattices nexp . However, there is still a lack of the studies for NExP on networks, as well as the analytical understanding of NExP process. The main aim of the study is to make progress in filling this gap by developing the exact mean-field solutions to NExP based on generating function methods in order to address the resilience of complex networks. We derive mean-field solutions to NExP on random networks, leading to an analytical insight to the effect of the no-exclaves rule and the anatomy of novel phase transitions displayed by the NExP process.
@@
The paper is organized as follows. In Sec. II, we introduce the no-exclaves percolation model and explain how to apply no-exclave rules. We also introduce the concept of the giant NExP component as the order parameter. Next, we derive an analytical theory by using generating function formalism, to compute the size of giant NExP component and the average NExP component size in random networks (Sec. III). Based on the theory, we identify and analyze the location of percolation transitions. Finally, we apply our theory to random regular networks, Erdős-Rényi networks, and real-world networks from empirical data (Sec. IV). The summary and discussions are presented in Sec. V.
@@ 2 No-Exclaves Percolation Model
In this section, we present the no-exclaves percolation on networks, which was originally studied for lattices nexp . The NExP process is defined as follows. Initially, all nodes are in an empty state in a given network. With the occupation probability $q$, we set each node to be occupied, randomly and independently. In terms of the network robustness, $1-q$ corresponds to failure probability. So far, the procedures are the same as the usual site percolation problem. Next, we identify the exclaves, which are the components composed of unoccupied nodes that are completely surrounded by the occupied nodes, except for the largest such one. Finally, the exclave gets merged into a bigger component comprising itself and all the neighboring occupied components. Note that during this process multiple connected components, and even multiple exclaves, can merge into a single, larger component. We refer to the resulting combined components as the NExP components. We then measure the size $\Phi$ of the largest NExP component as an order parameter and use occupation probability $q$ as a control parameter. The size $\Phi$ is equivalent to the probability that a randomly selected node belongs to the largest NExP component.
@@
Figure 1 shows an example of no-exclaves percolation. Each node in a network is occupied with occupation probability $q$ [marked blue in Fig. 1(a)] or empty with probability $1-q$ [marked white in Fig. 1(a)]. Then, we identify the exclaves red-shaded in Fig. 1(b), the components of connected empty nodes that are completely surrounded by occupied nodes. Practically, if all links leading out from the components composed of empty nodes connect to occupied nodes, the components correspond to exclaves. Note that the “giant” connected component composed of empty nodes is not an exclave. Next, the exclave components merge with the neighboring occupied components, becoming “virtually”-occupied themselves as shown in Fig. 1(c). After agglomerating all the exclaves and surrounding occupied components, we identify the NExP component, the connected component of both real- and virtually-occupied nodes [Fig. 1(d)]. The largest NExP component in (d) is in general greater than the largest connected component in (a).
@@
Figure 1: Schematic illustrations of NExP on a network. (a) Each node in the network is either occupied (blue) with a probability $q$, or empty with a probability $1-q$. (b) Exclaves (red), components of empty nodes surrounded by occupied nodes, are identified. (c) The exclaves merge with surrounding components. (d) The giant NExP component is identified, marked by the grey shaded area.
3 Generating Function Methods
@@
In this section, we derive the analytic solution of NExP on random graphs using generating function formalism, applicable to locally-treelike networks in the thermodynamic limit. We observe that according to the no-exclaves rule, there can be three categories of components in NExP on large networks: i) a giant NExP component, ii) small (non-giant) NExP components, and iii) a giant unoccupied component. A giant unoccupied component, which contains a finite fraction of nodes in the thermodynamic limit and extends to infinity, cannot be completely surrounded and thus is not an exclave. Note that we include isolated unoccupied components within the NExP component as there is no need to distinguish between them for our purposes.
@@
Based in this observation, our strategy for the analytical solutions of NExP on networks is the following: We first identify the giant unoccupied component that should be excluded when obtaining NExP components by using a standard technique for the site percolation newman . The major difference is that we here consider a giant “unoccupied” component rather than occupied one that has been normally focused on. Then we derive self-consistency equations for the probability that a node arrived by following a randomly chosen link belongs to small NExP component. The size of the giant NExP component, denoted as $\Phi$, is determined by the complementary probability to the sum of the giant unoccupied component and the small NExP components.
@@ 3.1 The size of giant unoccupied component
As noted above, we first calculate the size $S_{0}$ of a giant unoccupied component in order to obtain the size $\Phi$ of the giant NExP component. The size $S_{0}$ can be obtained by following a conventional technique for the site percolation problem newman . It is important to note that since we calculated the size of the giant unoccupied component, we determine the giant component among “unoccupied” nodes, rather than among occupied nodes as in ordinary site percolation problems.
@@
where the excess degree stands for the degree for a node reached by following a randomly chosen link newman , and $z$ stands for the mean degree, $z=\langle k\rangle=\sum_{k}kP(k)$. Let $R(s)$ be the probability that a randomly chosen node belongs to a small (non-giant) unoccupied component with size $s$. Here an unoccupied component stands for a connected component composed of solely unoccupied nodes. Then we define a generating function of $R(s)$ as ${H}_{0}(x)=\sum_{s=0}^{\infty}R(s)x^{s}$. We also define a generating function ${H}_{1}(x)$ that generates the size of a small unoccupied component reached by following a randomly chosen link.
On locally tree-like networks, the size of the small unoccupied component to which node $i$ reached by following a randomly chosen link belongs is either zero if node $i$ is occupied (with probability $q$) or the sum of the sizes of the unoccupied components to which node $i$’s neighbors belong plus one for node $i$ itself if node $i$ is unoccupied (with probability $1-q$). Therefore, ${H}_{1}(x)$ and $H_{0}(x)$ satisfy the following self-consistency equations
@@
We define $w$ as the probability for a node arrived at following a randomly chosen link does not belong to a giant unoccupied component. According to the definition of $H_{1}(x)$, the value of $w$ is identical to $H_{1}(1)$, as given by Eq. (3)
@@
In addition, as given the definition ${H}_{0}(x)$, we have the total probability that a randomly chosen node belongs to a small unoccupied component as ${H}_{0}(1)=\sum_{s=0}^{\infty}R(s)$. Note that it is not necessarily normalized such that ${H}_{0}(1)=1$ because there may exist a giant unoccupied component that is not included in $R(s)$. The size of the giant unoccupied component, $S_{0}$, on a random network is equal to the probability that a randomly chosen node belongs to the giant unoccupied component. In conclusion, once we obtain $w$, the size $S_{0}$ of the giant unoccupied component can be obtained as
@@ 3.2 Self-consistency equations for NExP component sizes
In this section, we attempt to derive the self-consistency equations for NExP component sizes. Let $Q(\phi)$ be the probability that a randomly chosen node belongs to a finite (non-giant) NExP component with size $\phi$ excluding both the giant NExP component and the giant unoccupied component. Then we define the generating function of $Q(\phi)$, represented as $F_{0}(x)=\sum_{\phi=0}^{\infty}Q(\phi)x^{\phi}$. We also define the generating function ${F}_{1}(x)$ that generates the size of finite NExP components for a node reached by a randomly chosen link. Note that $Q(\phi)$ may not be normalized to be unity since it excludes both a giant NExP component and a giant unoccupied component.
We then attempt to derive self-consistency equations of ${F}_{0}(x)$ and ${F}_{1}(x)$ on a locally tree-like network. Assume that if we arrive at node $i$ which belongs to a small NExP component by following a randomly chosen link, the size $\phi$ of the small component including node $i$ is one for the node $i$ plus the sum of the sizes of the small components to which node $i$’s neighbors belong. A key point to consider is that the probability of belonging to a size $\phi$ component can vary depending on whether node $i$ is i) occupied or ii) unoccupied.
@@
ii) On the other hand, if node $i$ is unoccupied, every neighbor of node $i$ cannot be a part of a giant unoccupied component. If this were the case, node $i$ would naturally belong to a part of the giant unoccupied component. Therefore, the probability that a neighbor of an “unoccupied” node belongs to a small component with size $\phi$ is generated by ${F}_{1}(x)$
@@ 3.3 The size of giant NExP component
The size $\Phi$ of a giant NExP component on a random network is identical to the probability that a randomly chosen node belongs to the giant NExP component. The complementary probability of $\Phi$ is that a randomly chosen node either belongs to a small NExP component, with probability $F_{0}(1)$, or belongs to a giant unoccupied component, with probability $S_{0}$. The probability $F_{0}(1)$ can be easily expressed as
@@
In summary, our strategy for obtaining the size $\Phi$ of the giant NExP component has been as follows. We first determine the probability $w$ that a node arrived by following a link does belong to the giant unoccupied component by solving Eq. (5), and subsequently the size $S_{0}$ of the giant unoccupied component by Eq. (6). In addition, we obtain the probability $u$ that a node arrived by following a link belongs to a small NExP component by Eq. (9). Putting all together, we can obtain the size $\Phi$ of the giant NExP component by Eq. (3.3).

### v2

@@ Abstract
No-exclaves percolation (NExP) is a nonlocal percolation process in which the components are formed not only by the connected occupied nodes but also by the agglomeration of empty nodes completely surrounded by the occupied nodes. It has been studied in low dimensions, displaying such novel phenomena as the discontinuous transition to complete percolation. However, its characteristics in complex networks are still unexplored. In this paper, we study the NExP on random networks by developing mean-field solutions using the generating function formalism. Our theory allows us to determine the size of the giant no-exclaves component as well as the percolation threshold, which are in excellent agreements with Monte Carlo simulations on random networks and some real-world networks. We show that on random networks NExP exhibits three phases and two transitions between them: the phases are characterized by the presence or absence of not only the giant NExP component but also the giant unoccupied component, which is the giant connected component composed solely of unoccupied nodes. This work offers theoretical understanding on the anatomy of phase transitions in the NExP process.
@@ 1 Introduction
Understanding percolation of complex networks has an impact on our ability to predict and control the behaviors of complex systems as it can provide essential information regarding the connectivity and robustness of complex systems [1, 2, 3, 4]. Classical percolation models are based on the notion of connectivity in which a pair of nodes is considered a part of the same connected component if there exists at least one connected path between them [1]. This model has provided a simple yet powerful theoretical framework across various phenomena, including epidemic spreading [5, 6], traffic flows [7], the stability of power grid [8, 9, 10], and the resilience of networks [2, 11, 12]. There have also been numerous studies on the variants of standard percolation on networks that have modified the concept of connectivity such as $k$-core percolation [13], bootstrap percolation [14], $k$-selective percolation [15], and mutual percolation [16, 17, 18, 19].
@@
One of the most important applications of percolation theory in complex networks is the robustness of networks against random failures or attacks [2, 11, 12, 20, 21]. The robustness of networked systems is often measured by how the size of the largest connected component responds when some nodes in networks are failed [2, 21]. The premise underlying the measurement is that it is essential for nodes to maintain their connectivity in order to function properly [2, 16]. Additionally, when nodes become non-functioning due to failure or detachment from the largest component these nodes are assumed to be unable to recover permanently in most studies on the robustness of networks [2, 3]. In reality, however, damaged nodes can often recover with communications to properly functioning nodes that surround the damaged nodes [19, 22, 23, 24, 25]. For instance, temporary breakdown in real-world networks such as the brain and financial systems can recover as long as the damage is localized.
@@
The no-exclaves percolation (NExP) is a model incorporating such a recovery rule that the component of failed nodes that is completely surrounded by active (unfailed) nodes recovers to be unfailed and merges into the surrounding unfailed components [26]. The failed component that is surrounded by unfailed nodes is referred to as the “exclave” which is the terminology in political geography and means a part of a district geographically isolated from the mainland by surrounding alien territories. Our model does not allow for exclaves, hence the name “no-exclaves percolation,” or NExP for short. After applying the no-exclaves rule for the recovery, we identify connected components and assess the robustness of networks using the concept of the giant NExP component. The no-exclaves rule in our model is closely related to the no-enclaves percolation on two dimensions offering a theoretical explanation for a motor-driven collapse of cytoskeletal systems [27, 28, 29, 30]. Note that the no-enclaves rule does not readily apply to complex networks; the no-exclaves rule, however, is applicable to generic networks, while keeping the ingredient of nonlocality intact. Furthermore, the no-exclaves rule implies long distance communications between active nodes in some sense [31]. From this perspective, it might have potential applicability to quantum communication and quantum network issues as well [32, 33].
@@
It is known from numerical simulations that in low-dimensional systems the no-exclaves rule significantly affects the behavior of percolation transitions in spatial networks and Euclidean lattices [26]. However, there is still a lack of study for NExP on networks as well as the analytical understanding of NExP process. The main aim of the study is to make progress in filling this gap by developing the exact mean-field solutions to NExP based on the generating function method in order to address the resilience of complex networks. We derive mean-field solutions to NExP on random networks, leading to an analytical insight to the effect of the no-exclaves rule and the anatomy of novel phase transitions displayed by the NExP process.
@@
The paper is organized as follows. In Sec. II, we introduce the no-exclaves percolation model and explain how to apply no-exclave rules. We also introduce the concept of the giant NExP component as the order parameter. Next, we derive an analytical theory by using the generating function formalism, to compute the size of the giant NExP component and the average NExP component size in random networks (Sec. III). Based on the theory, we identify and analyze the location of percolation transitions. Finally, we apply our theory to random regular networks, Erdős-Rényi networks, and real-world networks from empirical data (Sec. IV). The summary and discussions are presented in Sec. V.
@@ 2 No-Exclaves Percolation Model
In this section, we present the no-exclaves percolation on networks, which was originally studied for lattices [26]. The NExP process is defined as follows. Initially, all nodes are in an empty state in a given network. With the occupation probability $q$, we set each node to be occupied, randomly and independently. In terms of the network robustness, $1-q$ corresponds to failure probability. So far, the procedures are the same as the usual site percolation problem. Next, we identify the exclaves, which are the components composed of unoccupied nodes that are completely surrounded by the occupied nodes, except for the largest such one. Finally, the exclave gets merged into a bigger component comprising itself and all the neighboring occupied components. Note that during this process multiple connected components, and even multiple exclaves, can merge into a single, larger component. We refer to the resulting combined components as the NExP components. We then measure the size $\Phi$ of the largest NExP component as an order parameter and use occupation probability $q$ as a control parameter. The size $\Phi$ is equivalent to the probability that a randomly selected node belongs to the largest NExP component.
@@
Figure 1 shows an example of the no-exclaves percolation. Each node in a network is occupied with occupation probability $q$ [marked blue in Fig. 1(a)] or empty with probability $1-q$ [marked white in Fig. 1(a)]. Then, we identify the exclaves red-shaded in Fig. 1(b), the components of connected empty nodes that are completely surrounded by occupied nodes. Practically, if all links leading out from the components composed of empty nodes connect to occupied nodes, the components correspond to exclaves. Note that the “giant” connected component composed of empty nodes is not an exclave. Next, the exclaves merge with the neighboring occupied components, becoming “virtually”-occupied themselves as shown in Fig. 1(c). After agglomerating all the exclaves and surrounding occupied components, we identify the NExP component, the connected component of both real- and virtually-occupied nodes [Fig. 1(d)]. The largest NExP component in (d) is in general greater than the largest connected component in (a).
@@
Figure 1: Schematic illustrations of NExP on a network. (a) Each node in the network is either occupied (blue) with a probability $q$, or empty with a probability $1-q$. (b) Exclaves (red), components of empty nodes surrounded by occupied nodes, are identified. (c) The exclaves merge with surrounding components. (d) The largest NExP component is identified, marked by the grey shaded area.
3 Analytic Solution of NExP
@@
In this section, we derive the analytic solution of NExP on random graphs using the generating function formalism, applicable to locally-treelike networks in the thermodynamic limit. We observe that according to the no-exclaves rule, there can be three categories of components in NExP on large networks: i) a giant NExP component, ii) small (non-giant) NExP components, and iii) a giant unoccupied component. A giant unoccupied component, which contains a finite fraction of nodes in the thermodynamic limit and extends to infinity, cannot be completely surrounded and thus is not an exclave. Note that we include isolated unoccupied components within the NExP component as there is no need to distinguish between them for our purposes.
@@
Based in this observation, our strategy for the analytical solutions of NExP on networks is the following: We first identify the giant unoccupied component that should be excluded when obtaining NExP components by using a standard technique for the site percolation [34]. The major difference is that we here consider a giant “unoccupied” component rather than occupied one that has been normally focused on. Then we derive self-consistency equations for the probability that a node arrived by following a randomly chosen link belongs to small NExP component. The size of the giant NExP component, denoted as $\Phi$, is determined by the complementary probability to the sum of the giant unoccupied component and the small NExP components.
@@ 3.1 The size of giant unoccupied component
As noted above, we first calculate the size $S_{0}$ of a giant unoccupied component in order to obtain the size $\Phi$ of the giant NExP component. The size $S_{0}$ can be obtained by following a conventional technique for the site percolation problem [34]. It is important to note that since we calculated the size of the giant unoccupied component, we determine the giant component among “unoccupied” nodes, rather than among occupied nodes as in ordinary site percolation problems.
@@
where the excess degree stands for the degree for a node reached by following a randomly chosen link [34], and $z$ stands for the mean degree, $z=\langle k\rangle=\sum_{k}kP(k)$. Let $R(s)$ be the probability that a randomly chosen node belongs to a small (non-giant) unoccupied component with size $s$. Here an unoccupied component stands for a connected component composed solely of unoccupied nodes. Then we define a generating function of $R(s)$ as ${H}_{0}(x)=\sum_{s=0}^{\infty}R(s)x^{s}$. We also define a generating function ${H}_{1}(x)$ that generates the size of a small unoccupied component reached by following a randomly chosen link.
On locally tree-like networks, the size of the small unoccupied component to which the node $i$ reached by following a randomly chosen link belongs is either zero if $i$ is occupied (with probability $q$) or the sum of the sizes of the unoccupied components to which $i$’s neighbors belong plus one (for $i$ itself) if $i$ is unoccupied (with probability $1-q$). Therefore, ${H}_{1}(x)$ and $H_{0}(x)$ satisfy the following self-consistency equations
@@
We define $w$ as the probability for a node arrived at following a randomly chosen link does not belong to a giant unoccupied component. According to the definition of $H_{1}(x)$, the value of $w$ is identical to $H_{1}(1)$ and given by Eq. (3)
@@
In addition, from the definition ${H}_{0}(x)$, we have the total probability that a randomly chosen node belongs to a small unoccupied component as ${H}_{0}(1)=\sum_{s=0}^{\infty}R(s)$. Note that it is not necessarily normalized such that ${H}_{0}(1)=1$ because there may exist a giant unoccupied component that is not included in $R(s)$. The size of the giant unoccupied component, $S_{0}$, on a random network is equal to the probability that a randomly chosen node belongs to the giant unoccupied component. In conclusion, once we obtain $w$, the size $S_{0}$ of the giant unoccupied component can be obtained as
@@ 3.2 Self-consistency equations for NExP component sizes
In this section, we derive the self-consistency equations for NExP component sizes. Let $Q(\phi)$ be the probability that a randomly chosen node belongs to a finite (non-giant) NExP component with size $\phi$ excluding both the giant NExP component and the giant unoccupied component. Then we define the generating function of $Q(\phi)$, represented as $F_{0}(x)=\sum_{\phi=0}^{\infty}Q(\phi)x^{\phi}$. We also define the generating function ${F}_{1}(x)$ that generates the size of finite NExP components for a node reached by a randomly chosen link. Note that $Q(\phi=1)$ may not be normalized to be unity since it excludes both the giant NExP component and the giant unoccupied component.
We then derive self-consistency equations of ${F}_{0}(x)$ and ${F}_{1}(x)$ on a locally tree-like network. Assume that if we arrive at node $i$ which belongs to a small NExP component by following a randomly chosen link, the size $\phi$ of the small component including the node $i$ is one plus the sum of the sizes of the small components to which $i$’s neighbors belong. A key point to consider is that the probability of belonging to a size-$\phi$ component can vary depending on whether node $i$ is i) occupied or ii) unoccupied.
@@
ii) On the other hand, if node $i$ is unoccupied, every neighbor of node $i$ cannot be a part of a giant unoccupied component. If it were, node $i$ would naturally belong to a part of the giant unoccupied component. Therefore, the probability that a neighbor of an “unoccupied” node belongs to a small component with size $\phi$ is generated by ${F}_{1}(x)$.
@@ 3.3 The size of giant NExP component
We are ready to obtain the giant NExP component size $\Phi$ on a random network which is identical to the probability that a randomly chosen node belongs to the giant NExP component. The complementary probability of $\Phi$ is that a randomly chosen node either belongs to a small NExP component, with probability $F_{0}(1)$, or belongs to a giant unoccupied component, with probability $S_{0}$. The probability $F_{0}(1)$ can be easily expressed as
@@
In summary, our strategy for obtaining the size $\Phi$ of the giant NExP component has been as follows. We first determine the probability $w$ that a node arrived by following a link does not belong to the giant unoccupied component by solving Eq. (5), and subsequently the size $S_{0}$ of the giant unoccupied component by Eq. (6). In addition, we obtain the probability $u$ that a node arrived by following a link belongs to a small NExP component by Eq. (9). Putting all together, we can obtain the size $\Phi$ of the giant NExP component by Eq. (3.3).